# Complexity of Weighted First-Order Model Counting in the Two-Variable Fragment with Counting Quantifiers: A Bound to Beat

### v1

@@ 5.1.1 Sentences with Counting Quantifiers
5.1.2 Sentences with the Linear Order Axiom
We also consider sentences from the language of $\textbf{C}^{2}$ extended with the linear order axiom, i.e., one of the known extensions of $\textbf{C}^{2}$ that is still domain-liftable (Tóth and Kuželka, 2023). Sentence $\Gamma_{4}$ encodes a graph similar to the Model A of the Barabási-Albert model (Albert and Barabási, 2002), an algorithm for generating random networks.
$\Gamma_{4}$ $=\forall x\;Eq(x,x)$
$\wedge(|Eq|=n)$
@@ 5.1.2 Sentences with the Linear Order Axiom
In a sense, the graph encoded by $\Gamma_{4}$ on $n$ vertices is gradually grown. We start by ordering the vertices using the linear order axiom. Then, a complete graph $K_{k+1}$ is formed on the first $k+1$ vertices. Afterwards, we start growing the graph by appending remaining vertices $i\in\{k+2,k+3,\ldots,n\}$ one at a time. When appending a vertex $i$, we introduce $k$ outgoing edges that can only connect to the vertices $\{1,2,\ldots,i-1\}$, i.e., all the new edges have a form $(i,j)$ where $j\in\{1,2,\ldots,i-1\}$. Ultimately, when counting the number of such graphs, we may not be interested in the same solutions differing by vertex ordering only, so we can divide the final number by $n!$. The ordering through the linear order axiom is, however, a very useful modelling construct—without it, modelling graphs such as the one above would likely not be possible in a domain-lifted way.
From now on, let us refer to graphs defined by the sentence $\Gamma_{4}$ as BA$(k)$. Figure 4 depicts the runtime for counting the graphs BA(3) on $n$ vertices using the old and the new encoding. The problems lead to 16 and 8 valid cells, respectively.
Figure 4: Runtime for counting BA(3) graphs
5.1.3 Closing Remarks on Performance
The experiments confirm that the new encoding is indeed more efficient, outperforming the old encoding on all tested instances. While that should not come as a surprise, since we have derived a bound provably better than the old one, it is also not completely obvious because FastWFOMC uses many algorithmic tricks that partially make up for the inefficiencies of the old encoding. However, not even our new bound allowed us to compute WFOMC within reasonable time for large domain sizes on all the tested problems.
Still, one should keep in mind that the alternative, that is solving WFOMC by propositionalization to WMC, scales extremely poorly as was repeatedly shown in the lifted inference literature (Meert et al., 2014). Thus, we must rely on algorithms operating at the lifted level. Our work extends the domains that can be efficiently handled by lifted algorithms, but more work is needed to extend the reach of lifted inference algorithms further.
@@ 5.2 Lifted Inference in Markov Logic Networks
Inference (and also learning) in MLNs is reducible to WFOMC (Van den Broeck et al., 2014). It is thus possible to efficiently perform exact lifted inference over an MLN using WFOMC as long as the network is encoded using a domain-liftable language. The reduction works as follows: For each $(w_{i},\alpha_{i}(\textbf{x}_{i}))\in\Phi_{\mathds{R}}$, introduce a new formula $\forall\textbf{x}_{i}:\xi_{i}(\textbf{x}_{i})\Leftrightarrow\alpha_{i}(\textbf{x}_{i})$, where $\xi_{i}$ is a fresh predicate, and set $w(\xi_{i})=\exp(w_{i}),\overline{w}(\xi_{i})=1$ and $w(Q)=\overline{w}(Q)=1$ for all other predicates $Q$. Formulas in $\Phi_{\infty}$ are added to the theory as additional constraints. Denoting the new theory by $\Gamma$ and a query by $\phi$, we can compute the inference as
@@
Figure 5 shows the inference results for $n=8$ and various weights $w$. For both graphs, the distribution starts as binomial for small $w$. Obviously, if we do not care about satisfying the smoker-friend property, random assignment of truth values to $Sm(t_{1}),Sm(t_{2}),\ldots,Sm(t_{n})$ will lead to a binomial distribution.
As we increase the weight, the distributions gradually reverse to the opposite scenario, a U-shaped distribution that prefers the extremes. That is also intuitive, as we mostly work with one connected component where everyone is everyone’s friend. However, a random $3$-regular graph appears to go through the uniform distribution whereas our BA model seems not to. That is due to the fact that the BA model is, by construction, always one connected component as opposed to a 3-regular graph that can also consist of two connected components on $n=8$ vertices.
@@ 6 Related Work
This work builds on a long stream of results from the area of lifted inference (Poole, 2003; de Salvo Braz et al., 2005; Jha et al., 2010; Gogate and Domingos, 2011; Taghipour et al., 2013; Braun and Möller, 2016; Dilkas and Belle, 2023). Particularly, we continue in the line of research into the task of weighted first-order model counting (Van den Broeck et al., 2011; Van den Broeck, 2011; Van den Broeck et al., 2014; Beame et al., 2015; Kazemi et al., 2016; Kuusisto and Lutz, 2018; Kuželka, 2021; van Bremen and Kuželka, 2021b; Tóth and Kuželka, 2023; Malhotra and Serafini, 2023; Malhotra et al., 2023).
To the best of our knowledge, there is no other literature available on the exact complexity of computing WFOMC for $\textbf{C}^{2}$ sentences. The closest resource is the one proving $\textbf{C}^{2}$ to be domain-liftable (Kuželka, 2021), which we directly build upon and, in some sense, extend. Another relevant resource, concerned with designing an efficient algorithm for computing WFOMC over $\textbf{FO}^{2}$, is van Bremen and Kuželka (2021a) whose FastWFOMC algorithm remains state-of-the-art and which can be used as a WFOMC oracle required to deal with cardinality constraints.
@@ 7 Conclusion
The best existing bound for time complexity of computing WFOMC over $\textbf{C}^{2}$ is polynomial in the domain size (Kuželka, 2021). However, as we point out, the degree of the polynomial is exponential in the parameter $k$ of the counting quantifiers. Using the new techniques presented in this paper, we reduce the dependency of the degree on $k$ to a quadratic one, thus achieving super-exponential speedup of the WFOMC runtime with respect to the counting parameter $k$.
The new encoding has the potential of improving any applications of WFOMC over $\textbf{C}^{2}$, or making some applications tractable in the practical sense. We support this statement further in an experimental section, where we provide runtime measurements for computing WFOMC of several $\textbf{C}^{2}$ sentences and sentences from a domain-liftable $\textbf{C}^{2}$ extension.
It remains an open question whether the complexity can be reduced even further. Thus, we only consider our new bound as a bound to beat and we certainly hope that someone will beat it in the future.
@@ Appendix A Bit Complexity
All of the bounds presented in our paper assume mathematical operations to take constant time, which is a common assumption when it comes to asymptotic complexity. However, in the context of WFOMC, it is quite common that the algorithm ends up working with extremely large values. Consider counting the number of undirected graphs without loops on $n$ vertices, i.e., computing WFOMC of a formula
@@
In Kuželka (2021), the author derives a bound on the number of bits required to represent each summand in the WFOMC computation, which turns out to be polynomial in $n$ and $\log M$, where $M$ is a bound on the number of bits required to represent the weights. Total number of such summands is then also polynomial in $n$. Thus, the polynomial bound for domain-liftability is not violated.

### v2

@@ 5.1.3 Counting $k$-regular Directed Graphs
5.1.4 Counting with the Linear Order Axiom
Next, let us consider a sentence from the language of $\textbf{C}^{2}$ extended with the linear order axiom. Sentence $\Gamma_{4}$ encodes a graph similar to the Model A of the Barabási-Albert model (Albert and Barabási, 2002), an algorithm for generating random networks:
$\Gamma_{4}$ $=\left(\forall x\;Eq(x,x)\right)\wedge(|Eq|=n)$
@@ 5.1.4 Counting with the Linear Order Axiom
In a sense, the graph encoded by $\Gamma_{4}$ on $n$ vertices is sequentially grown. We start by ordering the vertices using the linear order axiom. Then, a complete graph $K_{k+1}$ is formed on the first $k+1$ vertices. Afterward, we start growing the graph by appending remaining vertices $i\in\{k+2,k+3,\ldots,n\}$ one at a time. When appending a vertex $i$, we introduce $k$ outgoing edges that can only connect to the vertices $\{1,2,\ldots,i-1\}$, i.e., all the new edges have a form $(i,j)$ where $j\in\{1,2,\ldots,i-1\}$. Ultimately, when counting the number of such graphs, we may not be interested in the same solutions differing by vertex ordering only, so we can divide the final number by $n!$. The ordering through the linear order axiom is, however, a very useful modeling construct—without it, modeling graphs such as the one above would likely not be possible in a domain-lifted way.
From now on, let us refer to graphs defined by the sentence $\Gamma_{4}$ as BA$(k)$. Figure 4(a) depicts the runtime for counting the graphs BA(3) on $n$ vertices using the old and the new encoding. The problems lead to 16 and 8 valid cells, respectively.
(a) Counting BA(3) graphs
(b) Inference on an MLN over BA(3) network
Figure 4: Runtimes of tasks performed on BA(3)
5.2 Remarks on Performance
The experiments above confirm that the new encoding is indeed more efficient, outperforming the old encoding on all tested instances. While that should not come as a surprise, since we have derived a bound provably better than the old one, it is also not completely obvious because FastWFOMC uses many algorithmic tricks that partially make up for the inefficiencies of the old encoding. However, not even our new bound allowed us to compute WFOMC within reasonable time for large domain sizes on all the tested problems.
Still, one should keep in mind that the alternative, that is, solving WFOMC by propositionalization to WMC scales extremely poorly, as was repeatedly shown in the lifted inference literature (Meert et al., 2014). Thus, we must rely on algorithms operating at the lifted level. Our work extends the domains that can be efficiently handled by lifted algorithms, but more work is needed to extend the reach of lifted inference algorithms further.
@@ 5.3 Lifted Inference in Markov Logic Networks
Inference (and also learning) in MLNs is reducible to WFOMC (Van den Broeck et al., 2014). It is thus possible to efficiently perform exact lifted inference over an MLN using WFOMC as long as the network is encoded using a domain-liftable language.
The reduction works as follows: For each $(w_{i},\alpha_{i}(\textbf{x}_{i}))\in\Phi_{\mathds{R}}$, introduce a new formula $\forall\textbf{x}_{i}:\xi_{i}(\textbf{x}_{i})\Leftrightarrow\alpha_{i}(\textbf{x}_{i})$, where $\xi_{i}$ is a fresh predicate, and set $w(\xi_{i})=\exp(w_{i}),\overline{w}(\xi_{i})=1$ and $w(Q)=\overline{w}(Q)=1$ for all other predicates $Q$. Formulas in $\Phi_{\infty}$ are added to the theory as additional constraints. Denoting the new theory by $\Gamma$ and a query by $\phi$, we can compute the inference as
@@
Figure 5 shows the inference results for $n=8$ and various weights $w$. For both graphs, the distribution starts as binomial for small $w$. Obviously, if we do not care about satisfying the smoker-friend property, random assignment of truth values to $Sm(t_{1}),Sm(t_{2}),\ldots,Sm(t_{n})$ will lead to a binomial distribution. As we increase the weight, the distributions gradually reverse to the opposite scenario, a U-shaped distribution that prefers the extremes. That is also intuitive, as we mostly work with one connected component where everyone is everyone’s friend. However, a random $3$-regular graph appears to go through the uniform distribution whereas our BA model seems not to. That is due to the fact that the BA model is, by construction, always one connected component as opposed to a 3-regular graph that can also consist of two connected components on $n=8$ vertices. Runtimes of the inference task for several domain sizes can be inspected in Figure 4(b).
@@ 6 Related Work
This work builds on a long stream of results from the area of lifted inference (Poole, 2003; de Salvo Braz et al., 2005; Jha et al., 2010; Gogate and Domingos, 2011; Taghipour et al., 2013; Braun and Möller, 2016; Dilkas and Belle, 2023). Particularly, we continue in the line of research into the task of weighted first-order model counting (Van den Broeck et al., 2011; Van den Broeck, 2011; Van den Broeck et al., 2014; Beame et al., 2015; Kazemi et al., 2016; Kuusisto and Lutz, 2018; Kuželka, 2021; van Bremen and Kuželka, 2021b; Malhotra and Serafini, 2022; Tóth and Kuželka, 2023; Malhotra and Serafini, 2023; Malhotra et al., 2023).
To the best of our knowledge, there is no other literature available on the exact complexity of computing WFOMC for $\textbf{C}^{2}$ sentences. The closest resource is the one proving $\textbf{C}^{2}$ to be domain-liftable (Kuželka, 2021), which we directly build upon and, in some sense, extend. Besides that, Malhotra and Serafini (2022) later proposed a slightly different approach to dealing with counting quantifiers, although they did not analyze the method’s exact complexity either. However, as shown in the appendix, their techniques are also super-exponential in the counting parameters, not offering any speedup. Another relevant resource, concerned with designing an efficient algorithm for computing WFOMC over $\textbf{FO}^{2}$, is van Bremen and Kuželka (2021a) whose FastWFOMC algorithm remains state-of-the-art and it can be used as a WFOMC oracle required to deal with cardinality constraints.
@@ 7 Conclusion
The best existing bound for the time complexity of computing WFOMC over $\textbf{C}^{2}$ is polynomial in the domain size (Kuželka, 2021). However, as we point out, the polynomial’s degree is exponential in the parameter $k$ of the counting quantifiers. Using the new techniques presented in this paper, we reduce the dependency of the degree on $k$ to a quadratic one, thus achieving a super-exponential speedup of the WFOMC runtime with respect to the counting parameter $k$.
The new encoding can potentially improve any applications of WFOMC over $\textbf{C}^{2}$ or make some applications tractable in the practical sense. We support this statement further in the experimental section, where we provide runtime measurements for computing WFOMC of several $\textbf{C}^{2}$ sentences and sentences from a domain-liftable $\textbf{C}^{2}$ extension.
It remains an open question whether the complexity can be reduced even further. Thus, we only consider our new bound a bound to beat, and we certainly hope that someone will beat it in the future.
@@ Appendix A Bit Complexity
All of the bounds presented in this paper assume mathematical operations to take constant time, which is a common assumption when it comes to asymptotic complexity. However, in the context of WFOMC, it is quite common for the algorithm to end up working with extremely large values.
Consider counting the number of undirected graphs without loops on $n$ vertices, i.e., computing WFOMC of a formula
@@
In Kuželka (2021), the author derives a bound on the number of bits required to represent each summand in the WFOMC computation, which turns out to be polynomial in $n$ and $\log M$, where $M$ is a bound on the number of bits required to represent the weights. The total number of such summands is then also polynomial in $n$. Thus, the polynomial bound for domain-liftability is not violated.
@@
Appendix B Complexity of an Alternative Approach to Counting Quantifiers
In this work, we focus on and improve techniques proposed in Kuželka (2021) for dealing with counting quantifiers when computing WFOMC. A slightly different approach to counting in WFOMC was later proposed in Malhotra and Serafini (2022). For completeness, this section contains a proof sketch of their techniques also being super-exponential in the counting parameters, same as the approach from Kuželka (2021).
Let us consider a concrete example of computing WFOMC over a formula encoding $m$-regular graphs, [^11] We denote the degree by $m$ in this case to easily differentiate it from a vector $\mathbf{k}$ used in Malhotra and Serafini (2022). i.e.,
$\Gamma$ $=\left(\forall x\;\neg E(x,x)\right)\wedge\left(\forall x\forall y\;E(x,y)\Rightarrow E(y,x)\right)$
$\wedge\left(\forall x\exists^{=m}y\;E(x,y)\right).$
Following Theorem 3 in Malhotra and Serafini (2022), when counting over $\Gamma$, we first replace the counting subformula by an atom $A(x)$ with a fresh predicate $A/1$, i.e.,
$\Gamma^{\prime}$ $=\left(\forall x\;\neg E(x,x)\right)\wedge\left(\forall x\forall y\;E(x,y)\Rightarrow E(y,x)\right)$
$\wedge\left(\forall x\;A(x)\right)\wedge\left(\forall x\;A(x)\Leftrightarrow\left(\exists^{=m}y\;E(x,y)\right)\right).$
Second, we define additional formulas
$\Phi_{1}$ $=\bigwedge_{i=1}^{m}\forall x\forall y\;P_{i}(x)\Rightarrow\neg\left(\left(A(x)\vee B(x)\right)\Rightarrow f_{i}(x,y)\right)$
$\Phi_{2}$ $=\bigwedge_{1\leq i<j\leq m}\forall x\forall y\;f_{i}(x,y)\Rightarrow\neg f_{j}(x,y)$
$\Phi_{3}$ $=\bigwedge_{i=1}^{m}\forall x\forall y\;f_{i}(x,y)\Rightarrow E(x,y)$
$\Phi_{4}$ $=\forall x\;B(x)\Rightarrow\neg A(x)$
$\Phi_{5}$ $=\forall x\forall y\;M(x,y)\Leftrightarrow\left(\left(A(x)\vee B(x)\right)\wedge E(x,y)\right)$
$\Phi_{6}$ $=|A|+|B|=|f_{1}|=|f_{2}|=\ldots=|f_{m}|=\frac{|M|}{m},$
where $B,M,f_{i}$’s and $P_{i}$’s are fresh predicates.
Third, we eliminate the counting sentence in $\Gamma^{\prime}$ using the new formulas as
$\Gamma^{\prime\prime}$ $=\left(\forall x\;\neg E(x,x)\right)\wedge\left(\forall x\forall y\;E(x,y)\Rightarrow E(y,x)\right)$
$\wedge\left(\forall x\;A(x)\right)\wedge\bigwedge_{i=1}^{5}\Phi_{i}.$
Finally, the counting problem can be posed as
$\textup{WFOMC}(\Gamma,n,w,\overline{w})=\sum_{(\mathbf{k},\mathbf{h})\models\Phi_{6}}\binom{n}{\mathbf{k}}\;\mathcal{T}_{m}(\Gamma^{\prime\prime},\mathbf{k},\mathbf{h},w,\overline{w}),$ (29)
where $\binom{n}{\mathbf{k}}$ is the multinomial coefficient, $\mathcal{T}_{m}(\Gamma^{\prime\prime},\mathbf{k},\mathbf{h},w,\overline{w})$ is some function whose complexity we can assume to be constant with respect to $n$, and the summation argument $(\mathbf{k},\mathbf{h})\models\Phi_{6}$ denotes all vector pairs $(\mathbf{k},\mathbf{h})$ filtered so that they satisfy the cardinality constraints in $\Phi_{6}$. We leave to Malhotra and Serafini (2022) what exactly it means for $(\mathbf{k},\mathbf{h})$ to satisfy $\Phi_{6}$. For our analysis, it suffices to say that vector $\mathbf{k}$ orders cells of $\Gamma^{\prime\prime}$ and $\mathbf{k}_{i}$ specifies how many constants from the domain satisfy the $i$-th cell in an interpretation. Note that, in our example, the filtering will not reduce the number of instances of vector $\mathbf{k}$ that needs to be summed over. Hence, the time complexity of evaluating Equation 29 is at least exponential in the dimensionality of $\mathbf{k}$.
Let us now inspect the structure of cells of $\Gamma^{\prime\prime}$ which will allow us to compute their total number, i.e., it will allow us to exactly determine the dimensionality of vector $\mathbf{k}$. All cells must contain $\neg E(x,x)$ and $A(x)$ due to the non-counting conjuncts in $\Gamma^{\prime}$. Due to $\Phi_{4}$, cells must further contain $\neg B(x)$ and, by $\Phi_{5}$, $\neg M(x,x)$ as well. Furthermore, due to $\Phi_{3}$, the cells must also contain $\neg f_{i}(x,x)$ for all $i$. Atoms enforced so far cause the consequents of all implications in $\Phi_{1}$ to be satisfied, hence, we are free to choose the truth values of all $P_{i}(x)$ atoms. Therefore, we have $2^{m}$ cells and the dimensionality of $\mathbf{k}$ is exponential in the counting parameter $m$. Overall, we conclude that the approach from Malhotra and Serafini (2022) for computing WFOMC over $\textbf{C}^{2}$ is, in fact, also super-exponential in the counting parameters.